Problem 1

Tallahassee Clinic projected the following budget information for 2018:

Total FFS Visit Volume

90,000 visits

Payer Mix:

     Blue Cross

40%

     Celtic Insurance Company

60%

Reimbursement Rates:

     Blue Cross

$25 per visit

     Celtic Insurance Company

$20 per visit

Variable Costs – Resource Inputs:

      Labor

48,000 total hours

      Supplies

100,000 total units

Variable Costs – Input Prices:

       Labor

$25 per hour

       Supplies

$1.50 per unit

Fixed Costs (overhead, plant, and equipment)

$500,000

Construct Tallahassee Clinic’s static operating budget for 2018. (See Exhibit 8.3, page 283. Note that there are four components that need to be included: Volume Assumptions, Revenue Assumptions, Cost Assumptions, and the Pro Forma Profit and Loss or P&L projected Statement.)

Problem 2

Refer to Problem 1 above. Tallahassee Clinic’s actual results for 2018 are shown in the table below:

Total FFS Visit Volume

100,000 visits

Payer Mix:

     Blue Cross

40%

     Celtic Insurance Company

60%

Reimbursement Rates:

     Blue Cross

$28 per visit

     Celtic Insurance Company

$18 per visit

Variable Costs – Resource Inputs:

      Labor

50,000 total hours

      Supplies

150,000 total units

Variable Costs – Input Prices:

       Labor

$28 per hour

       Supplies

$1.50 per unit

Fixed Costs (overhead, plant, and equipment)

$500,000

a. Construct Tallahassee Clinic’s flexible budget for 2018 and actual operating results for 2018. (Hint: place the three budgets side by side. See Exhibits 8.4 and 8.5).

b. What is the profit variance?

c. Wat is the revenue variance?

d. What is the cost variance?

e. Focus on the revenue side. What is the volume variance?

f. Focus on the revenue side. What is the price variance?

g. Focus on the cost side. What is the volume variance?

h. Focus on the cost side. What is the management variance?

3393000 what the volume level is On ner, the $100 009 m

Resats S Capitated visits rrs Spo.000 $1,00,000 $960,000 5 s4800 6o0000 640,000 Conelbutlion margin $630,000 400,000 5 24,000 5 130.000( too.oo)253.000) Conducting the Variance Analysis As esplainnd carler, varianoe analysis lnvolves comparing two amounts, with the variance being the difference betwoon the values For example, if at the beginning of the year, a honpital expected to make a profie of S2 million but aual resuils were a profie of $3 milion, the variance would be $1 million tendard valoe, in this case 52 million of profits, ja the proée

I ONLY NEED THE ANSWERS TO PROBLEM 2 BUT IT REFERENCED PROBLEM 1
Answer #1

(a) statement showing flexible budget and its comparison with actual

Static budget for 90000 visits flexible budget (at standard cost ) per unit flexible budget (at standard cost ) for 100000 visits Actual for 100000 visits variance
sales 4050000 45 $ 4500000 4600000 100000 A
Direct material 150000 1.5 per unit 225000 225000 -
Direct labor 100000 25 per hour 1250000 1400000 150000 A
Variable overheads - - - - -
Total variable cost 1150000 26.5 1475000 1625000 -
Contribution 1100000 18.5 1850000 2975000 -
Fixed cost 500000 5.56 500000 500000 -
Net profit 600000 12.94 1350000 2475000 1125000 A

(b) Profit variance

standard profit (budgeted visits - actual visits) = 12.94(90000-100000)=12.94*10000=129400 F

(c) revenue variance

price per visit * Budgeted visits- Price per visit * actual visits= (45*90000) - (46*100000) = 550000 F

(d) cost variance

actual cost = variable cost + fixed cost = 1625000+500000=2125000

budgeted cost = variable cost + fixed cost =1475000+500000=1975000

actual cost -budgeted cost = 150000 A

(e) volume variance of  revenue side = budgeted price (actual visits - budgeted visits ) = 45(100000-90000) = 450000

(f) price variance of revenue side = actual visits (actual price - budgeted price ) = 100000(46-45) = 100000 F

(g) volume variance from cost side =

  • variable overheads = 0
  • fixed overheads expenditure variance = budgeted fixed overheads - actual fixed overheads = 500000-500000 = 0
  • fixed overheads volume variance= standard fixed overheads - budgeted fixed overheads = 556000-500000 = 56000 F
  • fixed volume variance = fixed overheads expenditure variance + fixed overheads volume variance = 56000 F